1. (a) A body is falling free in a vacuum. The fall is necessarily related to the gravitational

acceleration g and the height h from which the body is dropped. Use dimensional

analysis to show that the velocity V of the falling body satisfies the relation

V / gh = constant.


1
Expert's answer
2021-04-15T06:38:29-0400

Dimension of V is "\frac{m}{sec}."

Dimension of g is "\frac{m}{sec^2}."

Dimension of h is "m."

Thus, dimension of "\sqrt{gh}" is "\sqrt{\frac{m}{sec^2}\cdot m}=\frac{m}{sec}=" dimension of V.


Therefore, the relation "\frac{V}{\sqrt{gh}}=constant."